Question 7 1 pts The average service time at the CSUS Financial Aid Office is 5 minutes and the average time between arrivals is 10.7. What is the hourly arrival rate? Question 8 1 pts The average service time at the CSUS Financial Aid Office is 7.7 minutes and the average time between arrivals is 10.1. On average, what percentage of time is the server idle? (use decimal format, eg, enter 50% as .50) Question 9 1 pts The average service time at the CSUS Financial Aid Offce is 5.2 minutes and the average time between arrivals is 9.2. What is the average number of students being serviced?
Answer #1

7. Lambda(arrival rate ) = 60 / 10.7 = 5.6 per hour

8. Lambda = 60 / 10.1 = 5.94

Mu = 60 / 7.7 = 7.79

Utilization = lambda / mu = 5.94/7.79 = 0.76 or 76% (busy)

Server idle = 100 - 76 = 24%

9. Average number of students serviced = service rate = 1 / service time

= 1 / 5.2 = 0.192 per minutes

Do give a positive rating if you think this answer helped